OpenStudy (anonymous):

What is the slope of the line that passes through the points (–6,1) and (4,–4)? A. -2 B. 2 C. -½ D. ½

OpenStudy (anonymous):

same formula.

OpenStudy (anonymous):

so 1-(-4)/-6-4 ?

OpenStudy (anonymous):

which is perfectly correct. :D

OpenStudy (anonymous):

would it be -1/2 or 1/2

OpenStudy (e.cociuba):

It would be -1/2

OpenStudy (anonymous):

1+4/-6-4 = 5/-10 = -5/10 = -1/2

OpenStudy (e.cociuba):

-5/10=-1/2